# PPA-Game: Characterizing and Learning Competitive Dynamics Among Online Content Creators

### v2

@@ Appendix C Omitted Proofs
$G(u^{\prime}-u^{\prime\prime})=-\frac{u^{\prime}-u^{\prime\prime}}{4+3\Gamma}+\frac{1}{3}\geq-\frac{1+\Gamma}{4+3\Gamma}+\frac{1}{3}>0\text{ and }G(u^{\prime}-u^{\prime\prime})=-\frac{u^{\prime}-u^{\prime\prime}}{4+3\Gamma}+\frac{1}{3}<\frac{1}{3}<\frac{1}{2}.$
(c) Finally, we could prove this lemma according to Theorem 1 in [Pradelski and Young, 2012]. Note that our Markov process $\mathcal{M}(\epsilon;\mathcal{G}^{\prime})$ follow exactly the one proposed in [Pradelski and Young, 2012] and the only modifications are to ensure that the payoff calculation for each player follows the game $\mathcal{G}^{\prime}$. By part (a) and (b) of our proof, the conditions of Theorem 1 in [Pradelski and Young, 2012] hold. As a result, since $\mathcal{G}^{\prime}$ have at least one PNE and the most efficient PNE is unique with probability $1$ according to Theorem 4.1, the only stochastically stable state is $Z^{*}$ in Lemma C.5. As a result, when $\epsilon\rightarrow 0$, $p_{Z^{*}}(\epsilon;\mathcal{G}^{\prime})\rightarrow 1$. Now the claim follows. ∎
@@
$H_{s}=\sum_{t\in w_{s}}\mathbb{I}[Z(t)=Z^{*}]\quad\text{and}\quad E_{2s}=\left\{\sum_{i=1}^{s}H_{i}>\frac{L_{s}}{2}\right\},L_{s}=\sum_{i=1}^{s}c_{2}i^{\eta}.$
For any $10<\alpha<$ 1, let $\epsilon$ be small enough that $p_{Z^{*}}(\epsilon;\mathcal{G}^{\prime})>1/(2(1-\alpha))$ according to Lemma C.5. Then there exists a constant $C(\epsilon;\mathcal{G}^{\prime})$, such that
@@
$\mathbb{P}\left[H_{s}\leq(1-\alpha)p_{Z^{*}}(\epsilon;\mathcal{G}^{\prime})c_{2}s^{\eta}\right]\leq\frac{C(\epsilon;\mathcal{G}^{\prime})}{\sqrt{p_{\mathcal{D}}(\epsilon;\mathcal{G}^{\prime})}}\exp\left(-\frac{\alpha^{2}p_{Z^{*}}(\epsilon;\mathcal{G}^{\prime})c_{2}s^{\eta}}{72A(\epsilon;\mathcal{G}^{\prime})}\right),$
@@
$E_{1}=\left\{\forall j\in[N],k\in[K],\left|\hat{\mu}_{j,k}-\mu_{k}\right|<\frac{\delta}{4K}\right\}.$
@@
(b) Consider the learning and exploitation phase. Under the clean event $E_{1}$, according to Theorem 4.1, we have that the most efficient Nash equilibrium $\bm{\pi}^{*}$ of game $\mathcal{G}^{\prime}$ is also a most efficient equilibrium of game $\mathcal{G}$. The solution is also unique with probability $1$. As a result, according to Lemma C.5, for any $10<\alpha<$ 1 there exists a small enough $\epsilon$ such that the state $Z^{*}$ in Lemma C.5 is the unique stationary distribution of $\mathcal{M}(\epsilon;\mathcal{G}^{\prime})$ and $p_{Z^{*}}(\epsilon;\mathcal{G}^{\prime})>1/(2(1-\alpha))$.
@@ C.6 Important Lemmas
Let $M$ be an ergodic Markov chain with state space $\mathcal{Z}$ and stationary distribution $p$. Let $A$ be its $1/8$-mixing time. Let $(Z_{1},\dots,Z_{t})$ denote a $t$-step random walk on $M$ starting from an initial distribution $\phi$ on $\mathcal{Z}$, i.e., $Z_{1}\sim\phi$. For every $i\in[t]$, let $f_{i}:\mathcal{Z}\rightarrow[0,1]$ be a weight function at step $i$ such that the expected weight $\mathbb{E}_{Z\sim p}[f_{i}(Z)]=\nu$ for all $i$. Define the total weight of the walk $(Z_{1},\dots,Z_{t})$ by $Y=\sum_{i=1}^{t}f_{i}(Z_{i})$. There exists some constant $c$ (which is independent of $\nu$ and $\alpha$) such that $\forall 0<\alpha<1$,

### v3

@@ Appendix C Omitted Proofs
[…]
(c) Finally, we could prove this lemma according to Theorem 1 in [Pradelski and Young, 2012]. Note that our Markov process $\mathcal{M}(\epsilon;\mathcal{G}^{\prime})$ follow exactly the one proposed in [Pradelski and Young, 2012] and the only modifications are to ensure that the payoff calculation for each player follows the game $\mathcal{G}^{\prime}$. By part (a) and (b) of our proof, the conditions of Theorem 1 in [Pradelski and Young, 2012] hold. As a result, since $\mathcal{G}^{\prime}$ have at least one PNE and the most efficient PNE is unique with probability 11 according to Theorem 4.1, the only stochastically stable state is $Z^{*}$ in Lemma C.5. As a result, when $\epsilon\rightarrow 0$, $p_{Z^{*}}(\epsilon;\mathcal{G}^{\prime})\rightarrow 1$. Now the claim follows. ∎
@@
[…]
For any $0<\alpha<1$, let $\epsilon$ be small enough that $p_{Z^{*}}(\epsilon;\mathcal{G}^{\prime})>1/(2(1-\alpha))$ according to Lemma C.5. Then there exists a constant $C(\epsilon;\mathcal{G}^{\prime})$, such that
@@
[…]
@@
[…]
@@
(b) Consider the learning and exploitation phase. Under the clean event $E_{1}$, according to Theorem 4.1, we have that the most efficient Nash equilibrium $\bm{\pi}^{*}$ of game $\mathcal{G}^{\prime}$ is also a most efficient equilibrium of game $\mathcal{G}$. The solution is also unique with probability 11. As a result, according to Lemma C.5, for any $0<\alpha<1$ there exists a small enough $\epsilon$ such that the state $Z^{*}$ in Lemma C.5 is the unique stationary distribution of $\mathcal{M}(\epsilon;\mathcal{G}^{\prime})$ and $p_{Z^{*}}(\epsilon;\mathcal{G}^{\prime})>1/(2(1-\alpha))$.
@@ C.6 Important Lemmas
Let $M$ be an ergodic Markov chain with state space $\mathcal{Z}$ and stationary distribution pp. Let AA be its $1/8$-mixing time. Let $(Z_{1},\dots,Z_{t})$ denote a $t$-step random walk on $M$ starting from an initial distribution $\phi$ on $\mathcal{Z}$, i.e., $Z_{1}\sim\phi$. For every $i\in[t]$, let $f_{i}:\mathcal{Z}\rightarrow[0,1]$ be a weight function at step $i$ such that the expected weight $\mathbb{E}_{Z\sim p}[f_{i}(Z)]=\nu$ for all ii. Define the total weight of the walk $(Z_{1},\dots,Z_{t})$ by $Y=\sum_{i=1}^{t}f_{i}(Z_{i})$. There exists some constant $c$ (which is independent of $\nu$ and $\alpha$) such that $\forall 0<\alpha<1$,